OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

I closed this. can you just give me the ansers?

OpenStudy (anonymous):

3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (dangerousjesse):

I'm just going to copy my answer from the previous question.. Solve the following system: {2 y-x = -6 3 x+y = 8 Choose an equation and a variable to solve for. In the second equation, look to solve for y: {2 y-x = -6 3 x+y = 8 Solve for y. Subtract 3 x from both sides: {2 y-x = -6 y = 8-3 x Perform a substitution. Substitute y = 8-3 x into the first equation: {2 (8-3 x)-x = -6 y = 8-3 x Expand the left hand side of the equation 2 (8-3 x)-x = -6. 2 (8-3 x)-x = (16-6 x)-x = 16-7 x: {16-7 x = -6 y = 8-3 x Choose an equation and a variable to solve for. In the first equation, look to solve for x: {16-7 x = -6 y = 8-3 x Isolate terms with x to the left hand side. Subtract 16 from both sides: {-7 x = -22 y = 8-3 x Solve for x. Divide both sides by -7: {x = 22/7 y = 8-3 x Perform a back substitution. Substitute x = 22/7 into the second equation

OpenStudy (dangerousjesse):

It's giving an example, not an answer.

OpenStudy (anonymous):

Answer?